An actor invests some money at 6​%, and ​$35000 more than three times three times the amount at 7%. The total annual interest earned from the investment is ​$38090. How much did he invest at each​ amount? Use the​ six-step method.
Answer by Boreal(15235)

x at 6%
3x+35,000 at 7%
.06x + (.07)(3x+35000)=38090
.06x+ .21x+2450=38090
0.27x=$35640
x=$132,000
3x+35000=$431,000
;
Check
132000@6%=$7920
431000@7%=$30170
Total is $38090